OpenStudy (anonymous):

A ring placed along y^2 + z^2 = 4, x = 0 carries a uniform charge of 5 uC/m. Find D at P(3,0,0).

OpenStudy (roadjester):

What is "D"?

OpenStudy (anonymous):

Electric Flux Density. Also using Coulombs laws. \[D = \epsilon _{0}*E\]

OpenStudy (roadjester):

Electric flux as in \[\huge \phi_e={q\over \epsilon_0}\]?

OpenStudy (anonymous):

We haven't used an equation with that notation. does q represent electric field intensity? If so then yes i do believe we are talking about the same thing. \[D/\epsilon _{0} = E <-- electricfield intenisity \]

OpenStudy (anonymous):

we have electric flux as \[\Psi = Q _{enc} (Charge enclosed) = \int\limits_{v}^{} \rho _{v} dv\] for example but D means electric flux density

OpenStudy (roadjester):

ok your notation is ENTIRELY different from my own textbook so I'm somewhat lost

OpenStudy (roadjester):

In my textbook coulomb's law is \[\huge F = k_e{q_1 q_2 \over r^2}\]q represents charge

OpenStudy (anonymous):

Alright we're still on the same page there!

OpenStudy (anonymous):

so long as \[k _{e} = \frac{ 1 }{ 4piepsilon _{0} }\] then I have electricfield intensity = \[\lim_{Q \rightarrow 0} \frac{ F }{ Q } \] Such that:\[E =\frac{ Q (r - r')}{ 4\pi \epsilon _{0} |r-r'|^3}\]

OpenStudy (roadjester):

okay and electric flux: \[\huge \phi_e = \int \vec E \dot{d\vec A}\]

OpenStudy (anonymous):

not quite... I have it as:

OpenStudy (anonymous):

\[\Psi = \int\limits_{}^{} D ds = Q _{enc}\]

OpenStudy (roadjester):

so you're looking for the electric field? If my E is your D then E represents the electric field

OpenStudy (anonymous):

Hmmm funny that they wouldn't use E but would use D to represent electric field. But okay. I am looking for electric flux density which is found by Electric flux Density = \[\int\limits_{region}^{} \frac{ p_{r}dr }{ 4piR^2}\]

OpenStudy (anonymous):

opps also multiply by the unit vector \[a _{r}\]

OpenStudy (roadjester):

just a minor favor, since the text is so small, can you enlarge it using \huge ?

OpenStudy (roadjester):

ok have you learned Gauss's law?

OpenStudy (anonymous):

okay i will in the future, we have but for this question it technically shouldn't be used yet. Gauss' law is the next section of questions that follows this one in the book titled 'gauss law and applications'.

OpenStudy (roadjester):

take a look at the picture i attached. I feel like that is essentially what you're dealing with. But with all of the differences between our texts, it's hard to make sure we're on the same page so I just uploaded that from my text

OpenStudy (roadjester):

hopefully it's what you need and makes SOME sense

OpenStudy (anonymous):

got the answer! It took a couple more steps but you got 90% of the way there. Thanks!

OpenStudy (roadjester):

sorry I couldn't help 100% :)

OpenStudy (anonymous):

You got me most of the way, more than i could have done!